Calculate the average kinetic energy of a gas at 300k in units of

J

kJ/mol

Cal/mol

eV

meV
Answer #1

Average Kinetic energy of the gas 1. Kang = 3 / 2 kot -23 K8 = Bolta manns constant (-38x 1ő I) Kang = 3X1.38X10 x 300% Tkan3:314 years ago the Ja can 8314 -23 Energy Karg = 621x10 J = 621x10x a Cal 8.314 = 149.38 xóts com 6.023x 10% Kany = 899,75 c